# CodeIF-Bench: Evaluating Instruction-Following Capabilities of Large Language Models in Interactive Code Generation

### v3

@@ 3.3. Benchmark Construction
3.3.1. VI Strategy Extraction
During software development, human code reviewers systematically assess code quality across multiple dimensions, subsequently providing actionable improvement suggestions. Developers then iteratively refine their code based on the review comment, forming a collaborative process known as code review (Li et al., 2022). The code review process exhibits similarities to the interactive process between programmers and LLM-generated code. Inspired by this analogy, we leverage real-world code review comments (Li et al., 2022) to extract instruction strategies, identifying common patterns for introducing follow-up requirements, aligning the distribution of human instructions in the real world. However, not all instruction types can be independently verified through test cases. As demonstrated in Figure 2, relatively subjective instructions (e.g., “What does # noqa mean?”) lack objective evaluation criteria, making them difficult to assess via automated testing. Furthermore, determining whether an instruction type can produce objective, verifiable outputs is non-trivial and often requires deep domain expertise. To address this, we employ manual annotation to filter verifiable instruction types serving as guidelines for generating diverse and actionable, verifiable instructions.
Table 2. The statistics of CodeIF-Bench. SA stands for StandAlone. Non-SA stands for repository-level code generation. Avg. L is the average lengths (tokens) of instructions calculated by GPT’s tokenizer.
@@ 3.3.1. VI Strategy Extraction
To extract diverse verifiable strategies, we first employ the KCenterGreedy (Sener and Savarese, 2018) algorithm, which has demonstrated effectiveness in selecting a well-distributed set of representative samples based on text embeddings, to select 500 candidate review comment examples. (In our experiment, 500 is far greater than the actual number of unique instructions.) Then, we prompt GPT-4o to generate strategies for each candidate comment, which we refer to as “High-level Requirement” to guide LLM to generate verifiable instructions. To guarantee the validity and verifiability of these strategies, we invited 2 experienced software engineering experts with at least 4 years of programming experience to check these strategies, by removing unverifiable instruction types, merging similar ones, and refining ambiguous requirements into high-quality VI strategies. All ambiguous samples were resolved through discussion and mutual agreement to ensure strict adherence to the above criteria.
3.3.2. Data Collection
To construct CodeIF-Bench, we first sample 124 programming tasks from two widely-used code generation benchmarks, where 50 tasks from MBPP (Austin et al., 2021) and 74 tasks from DevEval (Li et al., 2024a), encompassing both standalone functions and repository-level code generation. Specifically, for standalone function, we randomly sample 50 programming tasks from MBPP test dataset. The repo-level code generation tasks are sourced from DevEval, where the programming tasks are collected from 115 real-world projects across 10 domains covering different dependency types:
@@ 3.3.2. Data Collection
We select data based on the 10 repository classes and dependency types provided by DevEval, randomly selecting 2 programming tasks from each class, resulting in 20 programming tasks for each dependency type. Due to the limited number of cross-file programming tasks, resulting in only 14 data points. Ultimately, we obtain 74 repository-level programming tasks. Then we generate the VIs for each task based on the extracted instruction strategies, and we also build corresponding unit tests ($T$) for assessing whether the code follows the VI. Inspired by Ou et al. (2024), the instructions and tests are produced by LLMs, followed by human review and verification through execution. Specifically, the extracted VI strategies are applied to each task to guide the generation of verifiable instructions (7 for SA and 9 for Non-SA). Then we prompt GPT-4o to generate (VI, $T$) pairs based on the initial instruction and VI strategies. To further ensure data quality, we engaged two developers with over 4 years of Python programming experience to individually check and refine the data. They refined the dataset based on the following two criteria: 1. Ensuring instruction description accuracy—identifying vague descriptions or logical conflicts with the original programming task; 2. Validating test correctness—ensuring the test can verify instructions and introducing necessary dependencies to ensure proper execution of test functions. Through meticulous examination, they identified over 55% of the samples requiring refinement and achieved a Cohen’s Kappa score of 0.73. All ambiguous samples were resolved through discussion and mutual agreement to ensure strict adherence to the above criteria.
3.4. Dataset Overview
@@
Code Domain & Real World Scenario. CodeIF-Bench supports for both standalone function-level and repository-level code generation evaluation, where the programming tasks are derived from real-world projects (Li et al., 2024a) and crowd-sourced problems (Austin et al., 2021).
@@
Table 3. Benchmark comparison.
Benchmark Code Domain Real World Scenario Robust Metric Multi-Round
Code Generation Benchmarks
HumanEval (Chen et al., 2021)
MBPP (Austin et al., 2021)
CoderEval (Yu et al., 2024)
DevEval (Li et al., 2024a)
RepoEval (Zhang et al., 2023)
CanItEdit (Cassano et al., 2023)
NLP Instruction-Following Benchmarks
MT-Bench (Zheng et al., 2023)
Multi-IF (He et al., 2024)
InFo-Bench (Qin et al., 2024)
CodeIF (Yan et al., 2025)
CodeIF-Bench
Table 4. IA results in Static Conversation. Values with darker colors indicate higher performance.
L1 Claude-3.5-Sonnet 60.0 62.9 78.6 72.9 61.4 65.2 49.3 26.2 - - 59.9
GPT-4o 61.4 82.9 74.3 77.1 64.3 69.6 44.9 55.4 - - 66.4
DeepSeek-V3 52.9 74.3 74.3 67.1 45.7 68.1 37.7 55.4 - - 59.5
Qwen2.5-Coder-7B 45.7 68.6 71.4 60.0 38.6 65.2 33.3 41.5 - - 53.2
Qwen2.5-Coder-14B 48.6 70.0 70.0 60.0 38.6 65.2 30.4 43.1 - - 53.3
Qwen2.5-Coder-32B 51.4 75.7 75.7 75.7 47.1 68.1 39.1 46.2 - - 60.0
Qwen-Coder-Turbo 45.7 65.7 71.4 64.3 32.9 62.3 42.0 40.0 - - 53.2
L2 Claude-3.5-Sonnet 60.0 45.0 50.0 52.5 40.0 52.5 10.0 55.0 32.5 27.5 42.5
GPT-4o 35.0 35.0 40.0 37.5 12.5 17.5 5.0 20.0 15.0 12.5 23.0
DeepSeek-V3 42.5 52.5 62.5 40.0 32.5 37.5 10.0 45.0 40.0 32.5 39.5
Qwen2.5-Coder-7B 2.5 7.5 7.5 2.5 0.0 5.0 5.0 2.5 0.0 0.0 3.3
Qwen2.5-Coder-14B 17.5 15.0 20.0 12.5 7.5 7.5 7.5 10.0 5.0 2.5 10.5
Qwen2.5-Coder-32B 27.5 10.0 15.0 12.5 7.5 12.5 5.0 7.5 2.5 0.0 10.0
Qwen-Coder-Turbo 35.0 17.5 30.0 17.5 5.0 10.0 7.5 5.0 7.5 5.0 14.0
4. Evaluation
4.1. Research Questions
@@
  • RQ4: Exploration of IF Enhancement. How do existing fine-tuning and prompt-enhancement methods perform in enhancing instruction-following capabilities in interactive coding tasks?
4.2. Experimental Settings
@@
For Dynamic Conversation, the initial programming task serves as the instruction of the first round, and then we execute the corresponding tests to determine whether feedback rounds are needed to improve the code. Next, run all the tests corresponding to the verifiable instructions to select the instruction that does not satisfy the current round, and continue to execute the tests to determine whether feedback is required for the round. The maximum number of rounds is set to twice the total number of instructions (16 for L1 and 18 for L2), as this is the maximum number of interactions in the worst-case scenario.
@@ 4.3. Evaluated LLMs
  • Closed-Source LLM: We chose Claude-3.5-Sonnet (claude, 2023) and GPT-4o (gpt, 2024), powerful coding models in the field of code generation, and also introduced the closed-source models of the Qwen series, Qwen-Coder-Turbo (Hui et al., 2024).
  • Open-Source LLM: We select one of the open-source large-scale SOTA LLMs Deepseek-V3 (Zhu et al., 2024), and open-source small-scale SOTA code LLMs Qwen2.5-Coder-Instruct-7, 14, 32B (Hui et al., 2024).
Table 5. CA results in Static Conversation. Values with darker colors indicate higher performance.
@@
L1 Claude-3.5-Sonnet 60.0 54.3 55.7 48.6 35.7 35.7 19.3 - - 44.5
GPT-4o 61.4 68.6 51.4 44.3 35.7 22.9 10.4 - - 42.5
DeepSeek-V3 52.9 58.6 47.1 41.4 22.9 15.7 3.0 - - 34.8
Qwen2.5-Coder-7B 45.7 57.1 48.6 35.7 15.7 11.4 4.5 - - 31.5
Qwen2.5-Coder-14B 48.6 52.9 41.4 32.9 12.9 8.6 1.5 - - 28.6
Qwen2.5-Coder-32B 51.4 64.3 52.9 47.1 22.9 17.1 8.9 - - 38.1
Qwen-Coder-Turbo 45.7 48.6 41.4 34.3 11.4 5.7 3.0 - - 27.4
L2 Claude-3.5-Sonnet 60.0 37.5 30.0 25.0 10.0 5.0 0.0 0.0 0.0 18.6
GPT-4o 35.0 17.5 12.5 10.0 2.5 0.0 0.0 0.0 0.0 8.6
DeepSeek-V3 42.5 27.5 22.5 15.0 7.5 5.0 0.0 0.0 0.0 13.3
Qwen2.5-Coder-7B 2.5 2.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.6
Qwen2.5-Coder-14B 17.5 7.5 5.0 5.0 2.5 0.0 0.0 0.0 0.0 4.2
Qwen2.5-Coder-32B 27.5 2.5 2.5 2.5 2.5 2.5 0.0 0.0 0.0 4.4
Qwen-Coder-Turbo 30.0 5.0 2.5 2.5 2.5 0.0 0.0 0.0 0.0 4.7
5. Results and Analysis
5.1. RQ1: Performance in Static Conversation
To answer this research question, we evaluated the LLMs on CodeIF-Bench under the Static Conversation setting. The Instruction Accuracy (IA) results are presented in Table 4. In the L-1 task, all LLMs exhibited strong IF capabilities, with IA averages exceeding 50%, and Qwen2.5-Coder-32B-Instruct even achieves performance comparable to closed-source models like Claude and larger open-source LLM DeepSeek-V3. In L-2 task, performance varied substantially across LLMs. For example, Claude completed 60% of tasks in Turn-1, whereas one of the best open-source models, Qwen2.5-Coder-32B-Instruct, achieved less than 30%. Moreover, as the dialogue rounds progressed, Claude and DeepSeek-V3 maintained relatively robust instruction-following performance, but GPT-4o exhibited a notable decline from the 5th round onward, while open-source models with fewer parameters (e.g., the Qwen series) degraded sharply after the 2nd round. We hypothesize that longer contexts disrupt weaker LLMs, hindering their ability to follow current instructions. Another key observation is that these LLMs exhibited performance declines in similar rounds, such as the 5th, 7th, and 8th rounds in L-1 data and the 7th round in L-2. As mentioned in Section 3.1, the instruction order is pre-determined in the static conversation setting. These rounds predominantly involved involved three instruction categories: “Annotation Coverage”, “Code Standard”, and “Functionality Extension”. The LLM exhibited notably poor adherence to instructions in these categories, and we further investigated this phenomenon in RQ3 (Section 5.3).
The Conversation Accuracy (CA) results are shown in Table 5. Since the instruction for the last round (Turn 8 in L-1, Turn 10 in L-2) is “Functionality Extension”, which conflicts logically with other instructions, i.e., there is no code that passes both the “Functionality Extension” instruction and other instructions, we do not calculate the CA for Turn 8. On the L-1 dataset, all LLMs show varying degrees of performance degradation in CA as dialogue turns increase. Notably, nearly all models show substantial CA deterioration by the 7th round, and in L-2 (repository-level tasks), most models exhibit severe CA degradation starting from the second round, with near-complete breakdown in full-conversation instruction following by the 6th round. While most LLMs have relatively good ability to follow current instructions (as demonstrated in the IA results), LLM’s performance in following instructions throughout the entire conversation is not satisfactory. We attribute this to long-context interference, including the context of the task itself and the context interference caused by the dialogue, which accelerates forgetting of prior instructions.
Figure 3 presents the Instruction Forgetting Ratio (IFR) results of LLMs. Most LLMs demonstrate a gradual increase in IFR as dialogue rounds progress, indicating that the frequency of forgetting phenomena rises with the accumulation of conversational instructions. Moreover, LLMs’ performance on L-2 data exhibits significantly higher IFR than on L-1 data, implying that the contextual complexity of L-2 (the necessary repository context for code completion) plays a crucial role in weakening LLMs’ ability to retain instruction states. Notably, certain models, such as Claude, display fluctuating IFR trends over successive dialogue rounds, suggesting that internal reasoning mechanisms may enable implicit corrective behaviors without explicit feedback.
@@ 5.1. RQ1: Performance in Static Conversation
Figure 3. IFR results in Static Conversation.
RQ1 Summary: In Static Conversation without explicit external feedback, as the dialogue context expands, the LLM’s ability to follow current instructions (IA) and conversation instructions (CA) exhibits varying degrees of degradation influenced by model capabilities. While IA is further influenced by the instruction type, CA suffers from an additional decline due to instruction forgetting. where the Instruction Forgetting Rate (IFR) progressively increases as the conversation continues.

### v4

@@ 3.3 Benchmark Construction
3.3.1 VI Strategy Extraction
During software development, human code reviewers systematically assess code quality across multiple dimensions, subsequently providing actionable improvement suggestions. Developers then iteratively refine their code based on the review comment, forming a collaborative process known as code review [codereview]. The code review process exhibits similarities to the interactive process between programmers and LLM-generated code. Inspired by this analogy, we leverage real-world code review comments [codereview] to extract instruction strategies, identifying common patterns for introducing follow-up requirements, aligning the distribution of human instructions in the real world. However, not all instruction types can be independently verified through test cases. As demonstrated in Figure 2, relatively subjective instructions (e.g., “What does # noqa mean?”) lack objective evaluation criteria, making them difficult to assess via automated testing. Furthermore, determining whether an instruction type can produce objective, verifiable outputs is non-trivial and often requires deep domain expertise. To address this, we employ manual annotation to filter verifiable instruction types serving as guidelines for generating diverse and actionable, verifiable instructions.
To extract diverse verifiable strategies, we first employ the KCenterGreedy [kcenter] algorithm, which has demonstrated effectiveness in selecting a well-distributed set of representative samples based on text embeddings, to select 500 candidate review comment examples. (In our experiment, 500 is far greater than the actual number of unique instructions.) Then, we prompt GPT-4o to generate strategies for each candidate comment, which we refer to as “High-level Requirement” to guide LLM to generate verifiable instructions. To guarantee the validity and verifiability of these strategies, we invited 2 experienced software engineering experts with at least 4 years of programming experience to check these strategies, by removing unverifiable instruction types, merging similar ones, and refining ambiguous requirements into high-quality VI strategies. All ambiguous samples were resolved through discussion and mutual agreement to ensure strict adherence to the above criteria.
Table 2: The statistics of CodeIF-Bench. SA stands for StandAlone. Non-SA stands for repository-level code generation. Avg. L is the average lengths (tokens) of instructions calculated by GPT’s tokenizer
@@ 3.3.1 VI Strategy Extraction
3.3.2 Data Collection
To construct CodeIF-Bench, we first sample 124 programming tasks from two widely-used code generation benchmarks, where 50 tasks from MBPP [MBPP] and 74 tasks from DevEval [DevEval], encompassing both standalone functions and repository-level code generation. Specifically, for standalone function, we randomly sample 50 programming tasks from MBPP test dataset. The repo-level code generation tasks are sourced from DevEval, where the programming tasks are collected from 115 real-world projects across 10 domains covering different dependency types:
@@ 3.3.2 Data Collection
We select data based on the 10 repository classes and dependency types provided by DevEval, randomly selecting 2 programming tasks from each class, resulting in 20 programming tasks for each dependency type. Due to the limited number of cross-file programming tasks, resulting in only 14 data points. Ultimately, we obtain 74 repository-level programming tasks. Then we generate the VIs for each task based on the extracted instruction strategies, and we also build corresponding unit tests ($T$) for assessing whether the code follows the VI. Inspired by [ou2024inductive], the instructions and tests are produced by LLMs, followed by human review and verification through execution. Specifically, the extracted VI strategies are applied to each task to guide the generation of verifiable instructions (7 for SA and 9 for Non-SA). Then we prompt GPT-4o to generate (VI, $T$) pairs based on the initial instruction and VI strategies. To further ensure data quality, we engaged two developers with over 4 years of Python programming experience to individually check and refine the data. They refined the dataset based on the following two criteria: 1. Ensuring instruction description accuracy—identifying vague descriptions or logical conflicts with the original programming task; 2. Validating test correctness—ensuring the test can verify instructions and introducing necessary dependencies to ensure proper execution of test functions. Through meticulous examination, they identified over 55% of the samples requiring refinement and achieved a Cohen’s Kappa score of 0.73. All ambiguous samples were resolved through discussion and mutual agreement to ensure strict adherence to the above criteria.
3.4 Dataset Overview
Table 3: The comparison between popular benchmarks and CodeIF-Bench
Benchmark Code Domain Real World Scenario Robust Metric Multi-Round
Code Generation Benchmarks
HumanEval [HumanEval]
MBPP [MBPP]
CoderEval [CoderEval]
DevEval [DevEval]
RepoEval [RepoEval]
CanItEdit [CanItEdit]
NLP Instruction-Following Benchmarks
MT-Bench [MT-Bench]
Multi-IF [multiif]
InFo-Bench [InfoBench]
CodeIF [codeif]
CodeIF-Bench
@@ 3.4 Dataset Overview
Code Domain & Real World Scenario. CodeIF-Bench supports for both standalone function-level and repository-level code generation evaluation, where the programming tasks are derived from real-world projects [DevEval] and crowd-sourced problems [MBPP].
@@
4 Evaluation
4.1 Research Questions
@@
  • RQ4: Exploration of IF Enhancement. How do prompt-enhancement methods perform in enhancing instruction-following capabilities in interactive coding tasks?
4.2 Experimental Settings
@@
For Dynamic Conversation, the initial programming task serves as the instruction of the first round, and then we execute the tests in the instruction list (the initial state is all verifiable instructions) to select unfulfilled instructions with feedback in the next round. Next, we execute tests to verify whether the LLM-generated responses satisfy the current instruction. If the response fails, the current instruction is removed from the instruction list; otherwise, it is retained. This process continues until the instruction list becomes empty, the maximum number of interaction rounds is reached, or all instruction are successfully satisfied. In our experiments, the maximum number of interaction rounds is set to 9 for SA and 11 for Non-SA.
@@ 4.3 Evaluated LLMs
  • Closed-Source LLM: We chose Claude-3.5-Sonnet [claude] and GPT-4o [gpt4o], powerful coding models in the field of code generation.
  • Open-Source LLM: We select one of the open-source large-scale SOTA LLMs Deepseek-V3 [zhu2024deepseek], and open-source small-scale SOTA code LLMs Qwen2.5-Coder-Instruct-7, 14, 32B [qwencoder].
5 Results and Analysis
5.1 RQ1: Performance in Static Conversation
Table 4: IA results in Static Conversation. Values with darker colors indicate higher performance
L1 Claude-3.5-Sonnet 58.6 84.3 84.3 75.7 70.0 82.1 52.0 49.0 - - 69.5
GPT-4o 57.1 80.7 78.5 76.7 74.9 70.5 58.7 46.9 - - 64.0
DeepSeek-V3 60.0 80.0 84.3 77.4 55.7 71.8 29.1 54.9 - - 64.2
Qwen2.5-Coder-7B 44.3 65.7 62.9 58.6 35.7 66.0 11.7 41.5 - - 48.3
Qwen2.5-Coder-14B 48.6 72.9 70.0 64.3 40.0 72.0 17.4 44.5 - - 53.7
Qwen2.5-Coder-32B 51.4 75.7 75.7 68.6 50.0 69.0 17.5 41.7 - - 56.2
L2 Claude-3.5-Sonnet 55.0 60.0 62.5 52.5 37.5 25.0 7.5 50.0 30.0 37.5 41.8
GPT-4o 47.5 50.0 50.0 42.5 32.5 37.5 2.5 37.5 30.0 20.0 35.0
DeepSeek-V3 50.0 55.0 62.5 52.5 27.5 30.0 2.5 45.5 45.5 27.5 39.8
Qwen2.5-Coder-7B 27.5 32.5 42.5 22.5 22.5 20.0 2.5 20.0 12.5 12.5 21.5
Qwen2.5-Coder-14B 47.5 47.5 47.5 42.5 25.0 32.5 2.5 30.0 27.5 25.0 32.8
Qwen2.5-Coder-32B 42.5 37.5 35.0 27.5 12.5 15.0 2.5 15.0 7.5 2.5 19.8
Table 5: CA results in Static Conversation. Values with darker colors indicate higher performance
@@ 5.1 RQ1: Performance in Static Conversation
L1 Claude-3.5-Sonnet 58.6 79.8 80.3 80.2 78.3 76.2 68.0 - - 74.5
GPT-4o 57.1 80.7 78.5 76.7 74.9 70.5 58.7 - - 71.0
DeepSeek-V3 60.0 77.6 78.0 77.4 72.8 62.4 57.4 - - 69.4
Qwen2.5-Coder-7B 44.3 59.8 63.1 60.0 56.2 54.1 46.7 - - 54.9
Qwen2.5-Coder-14B 48.6 70.0 68.8 65.6 62.7 60.8 55.8 - - 61.7
Qwen2.5-Coder-32B 51.4 74.1 74.5 72.5 68.5 61.8 57.0 - - 65.7
L2 Claude-3.5-Sonnet 55.0 55.0 55.8 55.6 52.0 44.6 39.3 34.1 25.8 46.4
GPT-4o 47.5 51.2 48.3 46.9 35.0 36.3 31.1 29.7 27.8 39.3
DeepSeek-V3 50.0 51.2 54.2 51.2 35.0 32.1 29.3 31.6 33.6 40.9
Qwen2.5-Coder-7B 27.5 26.3 30.8 21.3 20.0 16.3 14.3 15.0 13.1 20.5
Qwen2.5-Coder-14B 47.5 45.0 44.2 43.1 30.0 32.5 27.9 26.3 25.8 35.8
Qwen2.5-Coder-32B 42.5 35.0 35.0 29.4 19.0 15.0 11.1 8.4 6.1 22.4
To answer this research question, we evaluated the LLMs on CodeIF-Bench under the Static Conversation setting. The Instruction Accuracy (IA) results are summarized in Table 4. In the L-1 task, most LLMs demonstrated strong abilities to follow users’ immediate instructions, with IA averages exceeding 50%. Notably, Qwen2.5-Coder-32B-Instruct achieved performance comparable to the large open-source model DeepSeek-V3. However, in the L-2 task, substantial performance disparities emerged across models. For example, Claude successfully completed 60% of the tasks in the first round, whereas Qwen2.5-Coder-7B-Instruct—one of the strongest open-source models—achieved a completion rate below 30%. Interestingly, Qwen2.5-Coder-14B-Instruct achieved first-round performance comparable to both the closed-source GPT-4o and the open-source SOTA model DeepSeek-V3. This finding suggests that carefully optimized training data and algorithms can enable smaller-parameter models to match larger open-source counterparts in single-round tasks. With increasing dialogue rounds, all models show reduced instruction-following capability, with the Qwen series experiencing a notably sharper decline than stronger models like Claude. This highlights the performance gap between open-source and closed-source models in interactions. These L-1 and L-2 results indicate that long-context management critically affects instruction following, and more capable LLMs better resist such long-context-induced degradation. It is worth noting that the performance differences among LLMs were most pronounced in Rounds 5, 7, and 8, likely due to the instruction types. This phenomenon is further examined in section 5.3.
Notably, Qwen2.5-Coder-14B-Instruct outperformed the 32B variant and was only marginally weaker than DeepSeek-V3 and GPT-4o. We observed that the 32B model often produced format errors during later dialogue turns—for instance, generating patch-style outputs while neglecting the initial instruction’s function requirement. These observations suggest that sustained instruction-following performance in multi-turn dialogues is primarily determined by a model’s exposure to multi-turn supervision rather than by parameter scale alone. Although Qwen2.5-Coder-32B-Instruct possesses greater capacity, its weaker performance relative to the 14B variant indicates that insufficient optimization for multi-turn instruction consistency can undermine its potential. This underscores that training strategies emphasizing dialogue-level instruction alignment are essential for maintaining stable adherence across turns.
@@
Figure 3: IFR results in Static Conversation
In the L-1 task, the Conversation Accuracy (CA) of all LLMs exhibits an initial rise followed by a gradual decline as dialogue rounds progress. This pattern suggests that LLMs can effectively follow both the user’s current instruction and prior instructions during the early interaction stages. The subsequent degradation can be attributed to two main factors. First, as the number of instructions and the volume of conversational history increase, models experience a reduced ability to follow the current instruction, as reflected by declining IA scores. Second, Figure 3 presents the instruction forgetting rate (IFR) results, showing that most models display a progressive increase in IFR with longer dialogues—indicating that LLMs tend to “forget” earlier instructions. This manifests in erroneous edits to previously completed instructions and poor control over the editing scope.
In the L-2 repository task, the introduction of additional repository context (averaging 9.8K tokens) led to a pronounced drop in CA metrics for nearly all models as the dialogue advanced. This was accompanied by further decreases in IA and substantial increases in IFR. These results reinforce that the repository context is a critical determinant of LLM performance during multi-turn interactions, exposing persistent limitations in instruction-following robustness under real-world development scenarios.
RQ1 Summary: Under the Static Conversation setting, both Instruction Accuracy (IA) and Conversation Accuracy (CA) decline as dialogue rounds increase: in L-1, CA rises initially but then drops, while IA decreases with accumulating instructions. Meanwhile, Instruction Forgetting Rate (IFR) steadily grows, indicating progressive forgetting and erroneous edits in multi-turn interactions. Adding repository context (L-2, about 9.8K tokens) amplifies these effects—CA and IA fall further, and IFR increases sharply.
5.2 RQ2: Performance in Dynamic Conversation